# HARMONY: Hidden Activation Representations and Model Output-Aware Uncertainty Estimation for Vision-Language Models

### v1

@@ Abstract
The growing deployment of Vision-Language Models (VLMs) in high-stakes applications such as autonomous driving and assistive technologies for visually impaired individuals necessitates reliable mechanisms to assess the trustworthiness of their generation. Uncertainty Estimation (UE) plays a central role in quantifying the reliability of model outputs and reducing unsafe generations via selective prediction. In this regard, most existing probability-based UE approaches rely on output probability distributions, aggregating token probabilities into a single uncertainty score using predefined functions such as length-normalization. Another line of research leverages model hidden representations and trains MLP-based models to predict uncertainty. However, these methods often fail to capture the complex multimodal relationships between semantic and textual tokens and struggle to identify biased probabilities often influenced by language priors. Motivated by these observations, we propose a novel UE framework, HARMONY, that jointly leverages fused multimodal information in model activations and the output distribution of the VLM to determine the reliability of responses. The key hypothesis of our work is that both the model’s internal belief in its visual understanding, captured by its hidden representations, and the produced token probabilities carry valuable reliability signals that can be jointly leveraged to improve UE performance, surpassing approaches that rely on only one of these components. Experimental results on three open-ended VQA benchmarks, A-OKVQA, VizWiz, and PathVQA, and three state-of-the-art VLMs, LLaVa-7b, LLaVA-13b and InstructBLIP demonstrate that our method consistently performs on par with or better than existing approaches, achieving up to 4% improvement in AUROC, and 6% in PRR, establishing new state of the art in uncertainty estimation for VLMs.
@@ 1 Introduction
Consider a visually impaired person querying a Vision-Language Model (VLM) with the question "What type of medicine is this?" unaware that the provided image may be blurry, unclear, occluded, or otherwise miss the information necessary to identify the drug. If model generates an answer without an uncertainty estimate, then the answer is not trustworthy for the person. If model generates an inaccurate answer with high confidence, or consistently outputs an incorrect answer, then acting on its generation can cause serious consequences for the person. One of the key research problems about the trustworthiness of VLMs is whether they can output reliable uncertainty estimate over the correctness of its generation or can they say ‘I don’t know’ instead of generating an incorrect answer for what they don’t know.
(a) Learnable UE scoring functions
(b) Challenges of the multimodal UE problem
Figure 1: Illustration of key challenges in UE problem for VLMs and how learnable scoring functions solve them. C1: Detect greater significance tokens. C2: learn how to aggregate token-level probability scores. C3: detect vision–text alignment especially for high-significance tokens. While LARS yaldiz2024not addresses the C1 and C2 using text and probability scores, and MSF dancette2023improving focuses on C3 by exploiting hidden states, our proposed method, HARMONY, jointly leverages text and probability scores to capture semantically significant tokens and token level uncertainty, and hidden representations to determine vision-text misalignment, thereby yielding a more reliable selective prediction estimate.
The UE problem is challenging due to auto-regressive nature of generation and multimodality. As shown in Figure 2, open-ended generations have multiple questions at the output, and some tokens may carry greater significance than the others and should be weighted more in UE estimation bakman2024mars . In addition, estimating uncertainty also involves aggregating the probabilities of individual tokens into a single UE score. Learning the aggregation function via heuristics such as mean srinivasan2024selective , product malinin2020uncertainty makes UE problem challenging due to factors such as length bias (shorter or longer responses affecting confidence score) malinin2020uncertainty , semantic bias (models favoring frequent phrases) bakman2024mars etc. Further, language prior is another challenge as VLMs are known for their tendency to overlook the evidence in the image, and over-rely on the language-priors avestimehr2025detecting .
@@
Many works have approached this problem via black-box formulations malinin2020uncertainty ; kuhn2023semantic ; tian2023just ; srinivasan2024selective ; avestimehr2025detecting . A key attraction of black-box methods is that they do not require training, and can work for proprietary-based models. In this regard, some works show that model’s consistency on its generation can be an indicator of its confidence khan2024consistency ; kuhn2023semantic . Others argue that self-prompting the model for its own generation can provide a better UE estimate tian2023just . Another line of work shows that evidence collection via asking relevant sub-questions can detect unreliable generation if the underlying VLM is well-calibrated, which in itself is a difficult condition to meet srinivasan2024selective . The other well-known formulation is white-box approach whitehead2022reliable ; dancette2023improving . This approach requires calibration datasets to train an auxiliary function. In this regard, prior works have shown that the hidden activation representations contain a multimodality reliability signal whitehead2022reliable ; dancette2023improving . They show its effectiveness by leveraging the representations of prompt, and answer to train an MLP-based reliability scoring function. The other work shows that training a transformer-like architecture on output probabilities can yield a good reliability score yaldiz2024not .
Complementing the findings of whitehead2022reliable and yaldiz2024not , we hypothesize that both model’s internal states carrying model’s internal understanding of the vision modality, and output probabilities capturing token-level uncertainty carry valuable reliability signal, and leveraging them both simultaneously can yield a better uncertainty estimate. Based on these insights, we present HARMONY (Hidden Activation Representations and Model Output-Aware Uncertainty Estimation for Vision-Language Models), a transformer architecture-based UE function that integrates generated text, their associated token probabilities, and the hidden representations of the model. Specifically, we employ VisualBERT li2019visualbert , a small-scale transformer with 113M parameters, which offers a relatively simple cost compared to training the original billion parameters VLM models.Through extensive experiments on three VQA benchmarks A-OKVQA, VizWiz, and PathVQA, and three frontier vision-language models (LLaVA-7B, LLaVA-13B, and InstructBLIP) and 8 existing UE baselines, we demonstrate that our method consistently performs on par with or better than existing black-box methods and learnable baselines, achieving up to 4% improvement in AUROC, 6% improvement in PRR, and up to 2.5% gain in the effective reliability metric, establishing new state-of-the-art performance in UE for VLMs.
@@ 2.1 Uncertainty Estimation
Given a question q, and an Image $\mathcal{I}$, a VLM model parameterized by $\theta$ generates an output response sequence $\textbf{s}=\{s_{1},s_{2},..,s_{k}\}$, where $k$ denotes the length of the sequence. The UE methods quantify the uncertainty for the model’s predicted sequence $s$ given the input context. A naive way of estimating uncertainty is to calculate the probability of a generated sequence,
@@
where $s_{<l}\overset{\triangle}{=}\{s_{1},s_{2},..,s_{l-1}\}$. Though there is no universally accepted definition of UE for LLMs and vision-language-models (VLMs) vashurin2025benchmarking , our work adopts a broadly accepted practical definition from previous works jiang2024graph ; yaldizdesign ; huang2024uncertainty , that is, for a given query q, image $\mathcal{I}$ and generated response s, an effective UE should assign a low uncertainty score (indicating higher confidence) if s is reliable in the given context. In tasks such as VQA evaluation benchmarks, reliability refers to the correctness of s with respect to the set of ground truth(s)yaldizdesign . Here, we present some of the state-of-the-art black-box UE methods.
Length-Normalized Scoring It is easy to note that the formulation given in 1 penalizes long sequences. Therefore, malinin2020uncertainty fixes the issue of length penalization in sequence probabilities by proposing the following proxy metric,
@@
First Token is another baseline that addresses the probability aggregation problem by leveraging only the confidence score of first token of the generated response $P(s_{0},|\textbf{q},\mathcal{I},\theta)$ .
@@ 3 Related Works
Output Consistency methods: uncertainty for these methods is estimated via examining the consistency of the generated output over multiple question rephrasings farquhar2024detecting ; khan2024consistency ; shah2019cycle or examining model confidence over relevant sub-questions srinivasan2024selective . The question rephrasings khan2024consistency or beam sampling based methods farquhar2024detecting are considered expensive due to multiple forward passes required of the large VLMs. Sub-question-based approaches srinivasan2024selective further add to the cost by requiring additional steps such as evidence collection, sub-question formulation, and relevance verification. Additionally, these methods assume that the VLM is well-calibrated, an assumption that does not always hold in practice.
Internal state examination methods: these works look at the model’s hidden activation representations dancette2023improving . Existing works whitehead2022reliable ; dancette2023improving exploit the representation vector of image, question and answer to predict the correctness of the response via an MLP-based learnable scoring function. While effective, these works require calibration datasets to train the function. Further, they train simple architectures such as learnable multi-layer perceptron (MLP)-based scoring functions to achieve the objective.
Token Probability methods: these methods use token probabilities assigned by the model at the output to predict the uncertainty malinin2020uncertainty ; kuhn2023semantic . In most cases, VLM-based UE methods frame open-ended visual question answering (VQA) tasks as multiple-choice problems khan2024consistency ; whitehead2022reliable . herefore, token probability methods remain relatively unexplored for generative VLMs. Some approaches leverage output probabilities and require calibration datasets for effective uncertainty estimation yaldiz2024not .
Our proposed method, HARMONY, integrates both internal state examination and token probability methods, combining their strengths to achieve a more robust UE framework.
@@ 4.1 Motivation
Semantic Significance and Inter-Token Dependencies: A well-calibrated model should exhibit a consistent relationship between its correctness and the probabilities it assigns to its predictions. In free-form generation, VLMs produce multiple tokens in an auto-regressive manner. Estimating uncertainty in this context involves aggregating the probabilities of individual tokens into a single uncertainty estimation score using a predefined scoring function. This makes UE inherently challenging due to factors such as length bias (shorter or longer responses affecting confidence score) malinin2020uncertainty , semantic bias (models favoring frequent phrases or syntactic structures) bakman2024mars that are often implicit, but significantly impact UE. Various functions proposed in the LLM literature aim to address different aspects of this aggregation process. For example, length-normalized scoring malinin2020uncertainty mitigates length bias, while semantic entropy kuhn2023semantic captures uncertainty across semantically similar responses. However, identifying an effective aggregation strategy through heuristics remains challenging due to inter-token dependencies and various inherent biases for the given vision and text context.
Language Priors: For VLMs, assessing token-level semantic significance alone may be insufficient for reliable uncertainty estimation. That is because VLMs are known for their tendency to overlook the evidence in the image, and over-rely on the language-priors. Consider the example question for which the model responds with ‘The man is looking up’. The uncertainty associated with the token ‘up’ should ideally reflect the model’s understanding of the visual scene. However, VLMs may assign a high probability to such tokens due to increasing confidence as the generation progresses, regardless of whether the visual input supports the claim. Therefore, the output probabilities can be biased towards language priors, and may not always be sufficient to yield a good reliability estimate. To address this challenge, we leverage hidden states. We discuss it further in Appendix Section A.
@@ 4.2 HARMONY
Let $f$ be the scoring function that takes four inputs: the question $\textbf{q}=(q_{1},q_{2},..,q_{K})$, the generated response $\textbf{s}=(s_{1},s_{2},..,s_{L})$, the token probabilities $\textbf{p}=(p_{1},p_{2},..,p_{L})$, and the model’s hidden-states $\textbf{H}=({\textbf{h}_{1},\textbf{h}_{2},..,\textbf{h}_{K},...,\textbf{h}_{K+L}})$ corresponding to q and s. Here, $p_{i}$ denotes the probability of generated sequence token $i$, and each hidden state $\textbf{h}_{i}\in R^{N}$ vector, where N represents the number of hidden units at a specific layer. It is worth-noting that the first $K$ vectors in H correspond to the question tokens, while the remaining $L$ represent the generated tokens. We use the hidden states of the tokens right after the visual tokens, as they inherently encode cross-modal interactions, capturing information transferred from visual tokens to text tokens. Note that our inputs consist of varied nature i.e, texts, probabilities which are real numbers, and hidden states which are high dimensional vectors requiring a structured approach to model their inter-dependencies in a sequential manner.
(a) VLM Inference
(b) Probability Token Embeddings
(c) HARMONY
Figure 2: An illustration of calibration data collection phase (left), probability token embedding design (middle) and the scoring function architecture (right). Subfigure 2(b) shows the orthogonal embedding vectors design of different probability quantiles. Subfigure 2(c) demonstrates how varied inputs are used to train a transformer like architecture, VisualBERT, to predict the reliability score.
@@
Learnable Task At the input, we have question, followed by generated tokens and their corresponding probabilities, further followed by hidden representations of the question and the generated tokens as shown in the Figure 2(c). Given these input, the task is to predict the correctness of the generation s for which we augment the VisualBERT model at the output via linear layer that gives a single logit output, $f(\textbf{q},\textbf{s},\textbf{H},\textbf{p})$. We employ binary cross-entropy loss,
$\mathcal{L}(f(\textbf{q},\textbf{s},\textbf{H},\textbf{p}),g)=-\left[g\log(f(\mathbf{q},\mathbf{s},\mathbf{H},\mathbf{p}))+(1-g)\log\left(1-f(\mathbf{q},\mathbf{s},\mathbf{H},\mathbf{p})\right)\right]$ (6)
where $g$ is the binary ground-truth label (of accuracy for generation s as an answer to q) used here as a target. Note that both VisualBERT model and the linear projection layer are fine-tuned on this reliability score prediction task.

### v2

@@ Abstract
Uncertainty Estimation (UE) plays a central role in quantifying the reliability of model outputs and reducing unsafe generations via selective prediction. In this regard, most existing probability-based UE approaches rely on predefined functions, aggregating token probabilities into a single UE score using heuristics such as length-normalization. However, these methods often fail to capture the complex relationships between generated tokens and struggle to identify biased probabilities often influenced by language priors. Another line of research uses hidden representations of the model and trains simple MLP architectures to predict uncertainty. However, such functions often lose the intricate inter-token dependencies. While prior works show that hidden representations encode multimodal alignment signals, our work demonstrates that how these signals are processed has a significant impact on the UE performance. To effectively leverage these signals to identify inter-token dependencies, and vision-text alignment, we propose HARMONY (Hidden Activation Representations and Model Output-Aware Uncertainty Estimation for Vision-Language Models), a novel UE framework that integrates generated tokens (’text’), model’s uncertainty score at the output (’MaxProb’), and its internal belief on the visual understanding of the image and the generated token (captured by ’hidden representations’) at token level via appropriate input mapping design and suitable architecture choice. Our experimental experiments across two open-ended VQA benchmarks (A-OKVQA, and VizWiz) and four state-of-the-art VLMs (LLaVA-7B, LLaVA-13B, InstructBLIP, and Qwen-VL) show that HARMONY consistently matches or surpasses existing approaches, achieving up to 5% improvement in AUROC and 9% in PRR.
Figure 1: Illustration of VLM inference and the proposed framework, HARMONY. VLMs as autoregressive model take image and question tokens as input and generate text tokens sequentially. During generation, certain tokens (e.g., “up" in the example) carry higher semantic relevance to the input, reflecting the inherently complex inter-token dependencies in VLM outputs. Moreover, some responses may be weakly grounded in the image and instead dominated by language priors, making UE particularly challenging. To address these issues, our proposed framework, HARMONY, jointly leverages the model’s hidden representations as a vision-text alignment signal together with the generated tokens and their sequential token-level probabilities, producing a single UE score for the generated answer.
@@ 1 Introduction
With the rise of Large-Language models, Vision-Language models (VLMs) have seen tremendous growth in recent years. Their remarkable performance in tasks such as visual reasoning, image captioning, and object detection has led to the growing deployment of these models in real-world applications, including assisting the Blind or Low Vision community ahmetovic2020recog, medical diagnosis sepehri2024mediconfusion, and robotic vision applications long2024robollm. However, VLMs often over-rely on language priors leng2024mitigating or co-occurring objects zhou2023analyzing in an image, which can lead to unreliable or misleading generations. In this regard, Uncertainty Estimation (UE) serves as a valuable approach for evaluating reliability. Unlike other hallucination detection methods that depend on external modules such as object detectors zhao2024mitigating, image generations for exploration augustin2025dash or image feature extractors deng2024seeing to infer vision entailment, UE derives reliability from the model’s own internal signals, offering a sophisticated means of assessing model’s uncertainty on its generations.
Since VLMs are autoregressive models, they may generate multiple tokens at the output. That is why, some of the key challenges of UE problem for VLMs are 1) capturing semantically significant tokens at the output generation, 2) learning the inter-token dependencies often reflected via the output distribution changes across tokens, and 3) identifying the vision-text alignment at token-level. We illustrate this via an example in Figure 1, where challenge 1) and 2) would entail identifying the relevant tokens to the given image and question, and learning how the UE distribution changes across tokens to formulate a UE score. Further, challenge 3 requires learning if those tokens are influenced by mere language priors (’model generates tokens based on text priors overlooking the image’) or are actually entailed in the image.
@@
Many works have approached the UE problem via black-box formulations malinin2020uncertainty; kuhn2023semantic; tian2023just; srinivasan2024selective; avestimehr2025detecting. A key attraction of these methods is that they do not require training, and can work for proprietary-based models. In this regard, some works show that model’s consistency can be an indicator of its confidence khan2024consistency; kuhn2023semantic. Others argue that self-prompting the model for its own generation can provide a better UE estimate tian2023just. Another line of work shows that evidence collection via asking relevant sub-questions can detect unreliable generation if the underlying VLM is well-calibrated, which in itself is a difficult condition to meet srinivasan2024selective.
The other well-known formulation is white-box approach whitehead2022reliable; dancette2023improving. This approach requires calibration datasets to train an auxiliary function. In this regard, prior works have shown that the hidden activation representations encode multimodal alignment signal whitehead2022reliable; dancette2023improving. They show its effectiveness by leveraging the representations of prompt, and answer to train an MLP-based reliability scoring function. However, such functions often lose the intricate inter-token dependencies. The other related work shows that training a transformer-like architecture on output probabilities can yield a good reliability score yaldiz2024not, however, how to encode multimodal alignment signal remains unexplored.
Given the inter-token dependencies and vision-text alignment challenge, and the findings of prior works whitehead2022reliable, yaldiz2024not, we hypothesize to leverage model’s internal states carrying model’s internal understanding of the vision modality, and output probabilities capturing token-level uncertainty such that output distribution shifts and internal belief on multimodal fusion are captured at token level. The key contributions of our works can be summarised as follows:
  • We propose a novel UE framework, HARMONY (Hidden Activation Representations and Model Output-Aware Uncertainty Estimation for Vision-Language Models), that leverages hidden activations, and output uncertainty scores to predict reliability score of the generated response.
  • Our experimental evaluations on 2 open-ended VQA benchmarks, 1) AOKVQA, a multimodal reasoning dataset, and 2) VizWiz, visually uncertain/unanswerable questions dataset collected by visually impaired people and four state-of-the-art VLMs (LLaVA-7B, LLaVA-13B, InstructBLIP, and Qwen-VL) show that HARMONY consistently matches or surpasses existing approaches, achieving up to 5% improvement in AUROC and 9% in PRR. We also include medical imaging dataset, PathVQA, evaluation comparison.
  • We also investigate selective prediction performance comparison of the learnable functions, aad further study the effectiveness of HARMONY for out-of-distribution generalization, calibration data sizes, and various input signal designs.
@@ 2.1 Uncertainty Estimation
Though there is no universally accepted definition of UE for LLMs and VLMs vashurin2025benchmarking, our work adopts a broadly accepted practical definition from previous works jiang2024graph; yaldizdesign; huang2024uncertainty, that is, for a given query q, image $\mathcal{I}$ and generated response s, an effective UE should assign a low uncertainty score (indicating higher confidence) if s is reliable in the given context. In tasks such as VQA evaluation benchmarks, reliability refers to the correctness of s with respect to the set of ground truth(s) yaldizdesign.
For VLMs, given a question q, and an Image $\mathcal{I}$, a model parameterized by $\theta$ generates an output response sequence $\textbf{s}=\{s_{1},s_{2},..,s_{k}\}$, where $k$ denotes the length of the sequence. The UE methods quantify the uncertainty for the model’s predicted sequence $s$ given the input context. A naive way of estimating uncertainty is to calculate the probability of a generated sequence,
@@
where $s_{<l}\overset{\triangle}{=}\{s_{1},s_{2},..,s_{l-1}\}$. It is easy to note that the formulation given in 1 penalizes long sequences. Below, we briefly describe some of the existing probability-based UE methods.
Length-Normalized Scoring malinin2020uncertainty fixes the issue of length penalization in sequence probabilities by proposing the following proxy metric,
@@
First Token zhao2024first is another baseline that addresses the probability aggregation problem by leveraging only the confidence score of first token of the generated response $P(s_{0},|\textbf{q},\mathcal{I},\theta)$ .
@@ 3 Related Works
Output Consistency methods: these methods estimate uncertainty via examining the consistency of the generated output over multiple question rephrasings farquhar2024detecting; khan2024consistency; shah2019cycle; zhang2024vl or examining model confidence over relevant sub-questions srinivasan2024selective. The question rephrasings khan2024consistency; zhang2024vl or beam sampling based methods farquhar2024detecting are considered expensive due to multiple forward passes required of the large VLMs. Sub-question-based approaches srinivasan2024selective further add to the cost by requiring additional steps such as evidence collection, sub-question formulation, and relevance verification. Additionally, these methods often assume that the VLM is well-calibrated srinivasan2024selective, an assumption that may not always hold in practice.
Internal state examination methods: these works look at the model’s hidden activation representations whitehead2022reliable; dancette2023improving. They have shown that internal representation of the autoregressive models carry significant reliability signal. They exploit the internal representation vector of input and answer, and train a simple scoring function, such as 2-layer MLP to predict the correctness of the response. While effective, these works require calibration datasets to train the function.
Token Probability methods: these methods use token probabilities at the output to predict the uncertainty by using pre-defined heuristic functions such as length-normalized scoring function malinin2020uncertainty, first token probability score zhao2024first etc. Some approaches leverage output probabilities and use learnable functions to aggregate probabilities for effective uncertainty estimation yaldiz2024not in open-ended question answering. In most cases, VLM-based UE methods frame open-ended visual question answering (VQA) tasks as multiple-choice problems khan2024consistency; whitehead2022reliable. Therefore, token probability methods remain relatively unexplored for generative VLMs.
Our proposed method, HARMONY, integrates both internal state examination and token probabilities, combining their strengths to achieve a more robust UE framework.
@@ 4.1 Motivation
Semantic Significance and Inter-Token Dependencies: In free-form generation, VLMs produce multiple tokens in an auto-regressive manner. Estimating uncertainty in this context involves aggregating the probabilities of individual tokens into a single uncertainty estimation score using a predefined scoring function. This makes UE inherently challenging due to factors such as length bias (shorter or longer responses affecting confidence score) malinin2020uncertainty, semantic bias (models favoring frequent phrases or syntactic structures) bakman2024mars that are often implicit, but significantly impact UE. Various functions proposed in the LLM literature aim to address different aspects of this aggregation process. For example, length-normalized scoring malinin2020uncertainty mitigates length bias, while semantic entropy kuhn2023semantic captures uncertainty across semantically similar responses. However, identifying an effective aggregation strategy through heuristics remains challenging due to inter-token dependencies and various inherent biases for the given vision and text context.
Language Priors: For VLMs, assessing token-level semantic significance and inter-token dependencies at the output alone can be insufficient for reliable uncertainty estimation. That is because VLMs are known for their tendency to overlook or misread the evidence in the image, and over-rely on the language-priors as shown in Figure 2. Consider the example question for which the model responds with ‘The man is looking up’. The uncertainty associated with the token ‘up’ should ideally reflect the model’s understanding of the visual scene. However, VLMs may assign a high probability to such tokens due to increasing confidence as the generation progresses, regardless of whether the visual input supports the claim. Therefore, we hypothesize that model’s internal belief at token level often captured by internal representations can help address language prior issue.
Figure 2: Illustration of language priors challenge where the LLaVa-7B model generates incorrect outputs with high confidence on visually uncertain (unanswerable) questions from the VizWiz dataset.
@@ 4.2 HARMONY
To address the inter-token dependencies, and language prior challenge, we aim to design a scoring function that looks at the semantics of the generation (’generated answer’: text), token-level uncertainty (’MaxProb’: a real valued number ), and model’s internal belief (’internal hidden representations’: large dimensional vector) of the model. In addition, we aim the function to learn inter-token dependencies, which requires to preserve the sequential nature of these signals as well. To achieve these objectives, we propose the following function, and input mappings.
Let $f$ be the scoring function that takes four inputs: the image conditioned question tokens $\textbf{q}=(q_{1},q_{2},..,q_{K})$, the generated response tokens $\textbf{s}=(s_{1},s_{2},..,s_{L})$, the token probabilities $\textbf{p}=(p_{1},p_{2},..,p_{L})$, and the model’s hidden-states $\textbf{H}=({\textbf{h}_{1},\textbf{h}_{2},..,\textbf{h}_{K},...,\textbf{h}_{K+L}})$ corresponding to q and s. Here, $p_{i}$ denotes the probability of generated sequence token $i$, and each hidden state $\textbf{h}_{i}\in R^{N}$ vector, where N represents the number of hidden units at a specific layer. It is worth-noting that the first $K$ vectors in H correspond to the question tokens, while the remaining $L$ represent the generated tokens. We use the hidden states of the tokens right after the visual tokens, as they inherently encode cross-modal fusion signal, capturing information transferred from visual tokens to text tokens. Note that our inputs consist of varied nature i.e, texts, probabilities which are real numbers, and hidden states which are high dimensional vectors requiring a structured approach to model their inter-dependencies in a sequential manner.
@@
Learnable Task At the input, we have question, followed by generated tokens and their corresponding probabilities, further followed by hidden representations of the question and the generated tokens as shown in the Figure 1. Given these input, the task is to predict the correctness of the generation s for which we augment the VisualBERT model at the output via linear layer that gives a single logit output, $f(\textbf{q},\textbf{s},\textbf{H},\textbf{p})$. We employ binary cross-entropy loss,
$\mathcal{L}\big(f(\mathbf{q},\mathbf{s},\mathbf{H},\mathbf{p}),g\big)$ $=-g\log f(\mathbf{q},\mathbf{s},\mathbf{H},\mathbf{p})$ (6)
$+(1-g)\log\big(1-f(\mathbf{q},\mathbf{s},\mathbf{H},\mathbf{p})\big)$ (7)
where $g$ is the binary ground-truth label (of accuracy for generation s as an answer to q) used here as a target. Note that both VisualBERT model and the linear projection layer are fine-tuned on this reliability score prediction task. It is worth mentioning that our proposed framework contributes at scoring function level, which can be integrated into other mixture-of-experts setup based baselines dancette2023improving.